Write sin(8x)- sin(2x) as a product of sines and cosines.
Answer #1

Formula used: sinca) – sinib) = 2 cos(Ate) sin(4-5) Answer : sin (8x) - Sincar) = 2 cos( 88 + ase ) sin(81-28) = dcos (103) s